Elliptic curves with large analytic order of �(E)

Andrzej Dąbrowski^{1} and Mariusz Wodzicki^{2}

1 Instytut Matematyki, Uniwersytet Szczeciński, ul. Wielkopolska 15, 70-451

Szczecin, Poland 2 Department of Mathematics, University of California, Berkeley, CA 94720-3840,

USA

To Yuri Ivanovich Manin on His Seventieth Birthday

Introduction

The L-series L(E, s) = ∑ _{n=1}^{∞} a_{n}n^{−s} of an elliptic curve E over ℚ converges for Res > 3/2. The Modularity Conjecture, settled by Wiles-Taylor-Diamond-Breuil-Conrad [BCDT], implies that L(E, s) analytically continues to an entire function and its leading term at s = 1 is described by the following long standing conjecture.

Conjecture 1 (Birch and Swinnerton-Dyer). L-function L(E, s) has a zero of order r = rank E(ℚ) at s = 1, and

lim_{s→1} \frac{L(E, s)}{(s − 1)^{r}} = \frac{c_{∞} (E)c_{fin}(E)R(E)|� (E)|}{|E(ℚ)_{tors} |^{2}} .

Here E(ℚ)_{tors} denotes the torsion subgroup of the group E(ℚ) of rational points of E, the fudge factor c_{fin} is the Tamagawa number of E, and R(E) is the regulator calculated with respect to the Néron-Tate height pairing. If ω is the real period of E, then c_{∞} = ω or 2ω, according to whether the group of real points E(ℝ) is connected or not.

Finally, � (E) denotes the Tate-Shafarevich group of E. The latter is formed by isomorphism classes of pairs (T, ϕ), where T is a smooth projective curve over ℚ of genus one which possesses a ℚ_{p}-rational point for every prime p (including p = ∞), and ϕ : E → Jac(T) is an isomorphism defined over ℚ. The Tate-Shafarevich group is very difficult to determine. It is known that the subgroup

� (E)[n] := {a ∈ � (E) | na = 0}

is finite for any n > 1 and it is conjectured that � (E) is always finite. In theory, the standard 2-descent method calculates the dimension of the 𝔽_{2}- vector space � (E)[2] (see [Cr_{1}], [S]). It is not clear in general how to exhibit the curves of genus 1 which represent elements of � (E) of order > 2 (see, however, [CFNS^{2}]).

It has been known for a long time that the order of � (E), provided the latter is always finite, can take arbitrarily large values. Cassels [C] was the first one to show this by proving that |� (E)[3]| can be arbitrarily large for a special family of elliptic curves with j-invariant zero. Only in 1987 it was finally established that there are any elliptic curves over ℚ for which the Tate-Shafarevich group is finite (Rubin [Ru], Kolyvagin [K], Kato). Ten years later Rohrlich [Ro] by combining results of [HL] and [K], proved that given a modular elliptic curve E over ℚ (hence any curve—according to [BCDT]), and a positive integer n, there exists a quadratic twist E_{d} of E such that � (E_{d}) is finite and |� (E_{d})[2]| n. This finally proved that � (E) can indeed be a group of arbitrarily large finite order.

Assuming the Birch and Swinnerton-Dyer Conjecture, Mai and Murty [M_{2}] showed that for the family of quadratic twists of any elliptic curve E, one has

\underline{lim}_{d} \frac{N(E_{d})^{\frac{1}{4} −ϵ}}{|� (E_{d})|} = 0.

Goldfeld and Szpiro [GS], and Mai and Murty [MM_{2}] (as reported by Rajan [R]), in the early 1990s proposed the following general conjecture:

Conjecture 2 (Goldfeld-Szpiro-Mai-Murty). For any ϵ > 0 we have^{3}

|� (E)| ≪ N(E)^{1/2+ϵ} .

Estimate (1) holds for the family of rank zero quadratic twists of any particular elliptic curve provided the Birch and Swinnerton-Dyer Conjecture holds for every member of that family.

The Birch and Swinnerton-Dyer Conjecture combined with the following consequence of the Generalised Lindelöf Hypothesis (see [GHP], p. 154)

lim_{d→∞} \frac{L^{(r_{d})} (E_{d}, 1)}{N(E_{d})^{ϵ}} = 0 (d square-zero),

where r_{d} denotes the rank of the group E_{d}(ℚ), and the following conjecture of Lang (see [L])

3In this article we adhere to the following notational convention. Let A(E) and B(E) be some quantities A(E) and B(E) dependent on a curve E belonging to a specified class 𝒞 of elliptic curves defined over ℚ. We say that A(E) ≪ B(E) if, for any K > 0, there exists N_{0} such that A(E) < KB(E) for all curves in 𝒞 with conductor N(E) > N_{0}. This is meaningful only if 𝒞 contains infinitely many nonisomorphic curves. If either A(E) or B(E) depend on some parameter ϵ, then the choice of N_{0} is allowed to depend on ϵ.

R(E) ≫ N(E)^{−ϵ} ,

easily imply that

|� (E_{d})| ≪ N(E_{d})^{1/4+ϵ} .

The following unconditional bounds

|� (E)| ≪ \{ \begin{matrix}N(E)^{79/120 +ϵ} & if j(E) = 0 \\ N(E)^{37/60 +ϵ} & if j(E) = 1728 \\ N(E)^{59/120 +ϵ} & otherwise, \end{matrix}

where j(E) denotes the j-invariant of E, are known for curves of rank zero with complex multiplication [GL].

In general, for elliptic curves satisfying the Birch and Swinnerton-Dyer Conjecture, Goldfeld and Szpiro [GS] show that the Goldfeld-Szpiro-Mai-Murty Conjecture is equivalent to the Szpiro Conjecture:

|Δ(E)| ≪ N(E)^{6+ϵ} ,

where Δ(E) denotes the discriminant of the minimal model of E. Masser proves in [Ma] that 6 in the exponent of (2) cannot be improved; in [We] de Weger conjectures that the exponent in (1) is also, in a certain sense, the best possible.

Conjecture 3 (de Weger). For any ϵ > 0 and any C > 0, there exists an elliptic curve over ℚ with

|� (E)| > CN(E)^{1/2−ϵ} .

He shows [We] that Conjecture 3 is a consequence of the following three conjectures: the Birch and Swinnerton-Dyer Conjecture for curves of rank zero, the Szpiro Conjecture, and the Riemann Hypothesis for Rankin-Selberg zeta functions associated to certain modular forms of weight \frac{3}{2}.

On the other hand, de Weger demonstrates that the following variant of Conjecture 3 which involves the minimal discriminant instead of the conductor, is a consequence of just the Birch and Swinnerton-Dyer Conjecture for elliptic curves with L(E, 1) ≠ 0.

Conjecture 4 (de Weger). For any ϵ > 0 and any C > 0, there exists an elliptic curve over ℚ with

|� (E)| > C|Δ(E)|^{1/12−ϵ} .

For the purpose of the present article the quantity

GS(E) := \frac{|� (E)|}{\sqrt{N(E)}}

will be referred to as the Goldfeld-Szpiro ratio of E. Eleven examples of elliptic curves with GS(E) 1 are given in [We], the largest value being 6.893.... Further forty seven examples with GS(E) 1 are produced by Nitaj [Ni], his largest value of GS(E) being 42.265. Note that curves of small conductor with GS(E) > 1 were already known from Cremona’s tables [Cr_{2}]. In all these examples GS(E) is calculated by using the formula for |� (E)| which is predicted by the Birch and Swinnerton-Dyer conjecture, see (4) below.

Let us say a few words about the order of the Tate-Shafarevich group for those curves when it is known. The results by Stein and his collaborators [GJPST, Thm. 4.4] imply that |� (E)| = 7^{2} for the curves denoted 546f2 and 858k2, respectively, in Cremona’s tables [Cr_{2}]. No other curve of rank zero and conductor less than 1000 has larger |� (E)| if the Birch and Swinnerton-Dyer conjecture holds for such curves. Gonzalez-Avilés demonstrated [GA, Thm. B], that formula (4) for the order of the Tate-Shafarevich group holds for all the quadratic twists

E_{d} : y^{2} = x^{3} + 21dx^{2} + 112d^{2}x

with L(E_{d}, 1) ≠ 0. The largest value of |� (E_{d})| for such curves, when d 2000, is |� (E_{1783} )| = 8^{2} (cf. [Le, Table I]).

Assuming the validity of the Birch and Swinnerton-Dyer conjecture, one can compute |� (E)| for an elliptic curve of rank zero E by evaluating L(E, 1) with sufficient accuracy. (In practice, this is possible only for curves with not too big conductors.) We shall be referring to this number as the analytic order of the Tate-Shafarevich group of E. In what follows |� (E)| will denote exclusively the analytic order of � (E).

It is rather surprising how small is the analytic order in all known examples: de Weger [We] produced one with |� (E)| = 224^{2}, Rose [Rs] produced another one with |� (E)| = 635^{2}; finally, Nitaj [Ni] found a curve with

|� (E)| = 1832^{2}

and that seems to be the largest known value prior to year 2002.

For the family of cubic twists considered by Zagier and Kramarz [ZK]

E′_{d} : x^{3} + y^{3} = d (d cubic-free),

the value of |� (E′_{d})| does not exceed 21^{2} for d 70000. In this case, the Birch and Swinnerton-Dyer, the Lang, and the Generalised Lindelöf conjectures imply that

|� (E′_{d})| ≪ N(E′_{d})^{1/3+ϵ} .

For quadratic twists of a given curve one can calculate the analytic order of the Tate-Shafarevich group by using a well known theorem of Waldspurger [W] in conjunction with purely combinatorial methods. The details for some curves with complex multiplication can be found in [Fr_{1}],[Fr_{2}],[Le],[N],[T]. Here we shall consider only one example, the family

E_{d} : y^{2} = x^{3} − d^{2}x (d 1 an odd square-free integer)

of so called congruent-number elliptic curves. Define the sequence a(d) by

∑_{n=1}^{∞} a(n)q^{n} := η(8z)η(16z)Θ(2z)

where

η(z) = q^{1/24} ∏_{n=1}^{∞} (1 − q^{n}),  Θ(z) = ∑_{n=−∞}^{∞} q^{n^{2}}  (q = e^{2πiz} ),

When curve E_{d} is of rank zero then, assuming as usual the Birch and Swinnerton-Dyer conjecture, we have (see [T]):

|� (E_{d})| = ( \frac{a(d)}{τ(d)}) ^{2}

where τ(d) denotes the number of divisors of d. (Coefficients a(d) can also be calculated using a formula of Ono [O].) Conjecturally, one expects that

|� (E_{d})| ≪ N(E_{d})^{1/4+ϵ} ,

hence the sequence of curves E_{d} (and, more generally, the family of quadratic twists of any curve) is not a likely candidate to produce curves with large Goldfeld-Szpiro ratio.

The primary aim of this article is to present the results of our search for curves with exceptionally large analytic orders of the Tate-Shafarevich group. We exhibit 134 examples of curves of rank zero with |� (E)| > 1832^{2} which was the largest previously known value for any explicit curve. For our record curve we have

|� (E)| = 63, 408^{2}.

For the reasons explicated in the last section, we focused on the family

E(n, p) : y^{2} = x(x + p)(x + p − 4 · 3^{2n+1} ),

and three families of isogeneous curves, for n and p being integers within the bounds 3 n 19 and 0 < |p| < 1000. Compared to the previously published results, in our work we faced dealing with curves of very big conductor. A big conductor translates into a very slow convergence rate of the approximation to L(E, 1). The main difficulty was to design a successful search strategy for curves with an exceptionally large Goldfeld-Szpiro ratio, (3), which is usually accompanied by a large value of the analytic order of the Tate-Shafarevich group.

Our explorations brought out also a number of unplanned discoveries: curves of rank zero with the value of L(E, 1) much smaller, or much bigger, than in any previously known example (see Tables 6 and 5 below). A particularly notable case involves a pair of non-isogeneous curves whose values of L(E, 1) coincide in their first 11 digits after the decimal!

Details of the computations, tables and related comments are contained in Sections 1 - 3. Further remarks on Conjecture 3 are the subject of Section 4.

The actual calculations were carried out by the second author in the Summer and early Fall 2002 on a variety of computers, almost all of them located in the Department of Mathematics in Berkeley. Supplemental computations were conducted also in 2003 and the Summer 2004.

The results were reported by M.W. at the conference Geometric Methods in Algebra and Number Theory which took place in December 2003 in Miami, and by A.D. at the Number Theory Seminar at the Max-Planck-Institut in October 2006; A.D. would like to thank the Department of Mathematics in Berkeley and the Max-Planck-Institut in Bonn, for their support and hospitality during his visits in 2006 when the revised version of this article was prepared; M.W. would like to thank the Institute of Matematics at the University of Szczecin for its hospitality during his visits there in the Summer 2002, when the project started, and in the Summer 2003. The second author was partially supported by NSF Grants DMS–9707965 and DMS–0503401.

1 Examples of elliptic curves with large |� (E)|

Consider the family

E(n, p) : y^{2} = x(x + p)(x + p − 4 · 3^{2n+1} ),

with (n, p) ∈ ℕ × ℤ and p ≠ 0, 4 · 3^{2n+1} . Any member of this family is isogeneous over ℚ to three other curves E_{i}(n, p) (i = 2, 3, 4):

E_{2}(n, p) : y^{2} = x^{3} + 4(2 · 3^{2n+1} − p)x^{2} + 16 · 3^{4n+2} x,  (1)
E_{3}(n, p) : y^{2} = x^{3} + 2(4 · 3^{2n+1} + p)x^{2} + (4 · 3^{2n+1} − p)^{2}x,  (2)

and

E_{4}(n, p) : y^{2} = x^{3} + 2(p − 8 · 3^{2n+1} )x^{2} + p^{2}x.  (3)

The L-series and ranks of isogeneous curves coincide, while the orders of E(ℚ)_{tors} and � (E), the real period, ω, and the Tamagawa number c_{fin} may differ. The curves being 2-isogeneous, the analytic orders of � (E_{i}) may differ from |� (E(n, p)| only by a power of 2.

All the examples we found where at least one of the four analytic orders of � (E(n, p) and � (E_{i}(n, p)) (i = 2, 3, 4) is greater or equal to 1000^{2} are listed in Table 1. Notation used: |� | = |� (E)| and |� _{i}| = |� (E_{i})|.

For a curve E of rank zero, we compute the analytic order of � (E), i.e., the quantity

|� (E)| = \frac{L(E, 1) · |E(ℚ)_{tors} |^{2}}{c_{∞} (E)c_{fin}(E)},

by using the following approximation to L(E, 1), cf. [Co]:

S_{m} = 2 ∑_{l=1}^{m} \frac{a_{l}}{l} e^{− \frac{2πl}{\sqrt{N}}} ,

which, for

m \frac{\sqrt{N}}{2π} ( 2 log2 + k log10 − log(1 − e^{−2π/\sqrt{N}} )) ,

differs from L(E, 1) by less than 10^{−k} .

It seems that the currently available techniques of n-descent for n = 3, 4, and 5 (cf. [CFNS^{2}], [MS^{2}], [Be], [F]), can be utilized to see that 60^{2} divides the actual order of � (E) for E = E_{3}(15, 12). On the other hand, the results of Kolyvagin and Kato could be used to prove that the actual order of � (E) divides |� (E)|. This would establish validity of the exact form of the Birch and Swinnerton-Dyer Conjecture in this case. The Birch and Swinnerton-Dyer conjecture is invariant under isogeny, hence this would establish validity of this conjecture for each of its three isogeneous relatives. In particular, this would show that � (E_{4}(15, 12)) is indeed a group of order 3840^{2}.

Table 1. Examples of elliptic curves E(n, p) (n 19; 0 < |p| 1000) with max(|� |, |� _{2}|, |� _{3}|, |� _{4}|) 1000^{2}.

\begin{matrix}(n, p) & N(n, p) & |� | & |� _{2}| & |� _{3}| & |� _{4}| \\ (11, −489) & 1473152464197864 & 680^{2} & 680^{2} & 1360^{2} & 680^{2} \\ (11, 163) & 1473152461647240 & 346^{2} & 1384^{2} & 173^{2} & 1384^{2} \\ (11, 301) & 5440722586421136 & 576^{2} & 1152^{2} & 576^{2} & 288^{2} \\ (11, 336) & 15816054028824 & 529^{2} & 1058^{2} & 529^{2} & 1058^{2} \\ (11, 865) & 15635299103673360 & 617^{2} & 1234^{2} & 617^{2} & 617^{2} \\ (12, −605) & 4473683858657640 & 1031^{2} & 1031^{2} & 1031^{2} & 2062^{2} \\ (12, −257) & 20904304573762872 & 1545^{2} & 1545^{2} & 3090^{2} & 3090^{2} \\ (12, −56) & 569377945555104 & 1049^{2} & 1049^{2} & 2098^{2} & 1049^{2} \\ (12, 22) & 143157883450560 & 416^{2} & 1664^{2} & 416^{2} & 1664^{2} \\ (12, 24) & 81339706505952 & 603^{2} & 1206^{2} & 603^{2} & 1206^{2} \\ (12, 63) & 63264216170568 & 554^{2} & 1108^{2} & 554^{2} & 1108^{2} \\ (12, 262) & 42622006206125760 & 468^{2} & 1872^{2} & 234^{2} & 1872^{2} \\ (12, 382) & 62143535763983040 & 648^{2} & 2592^{2} & 324^{2} & 2592^{2} \\ (12, 466) & 75808606453660608 & 1435^{2} & 5740^{2} & 1435^{2} & 5740^{2} \\ (12, 694) & 112899512607942336 & 576^{2} & 2304^{2} & 288^{2} & 2304^{2} \\ (12, 934) & 151942571712321216 & 512^{2} & 2048^{2} & 256^{2} & 2048^{2} \end{matrix}

(n, p)  N(n, p)  |� |  |� _{2}|  |� _{3}| |� _{4}|
(13, −672)  1281100377506040  389^{2} 1556^{2}  389^{2}  778^{2}
(13, −160)  915071698203240 1079^{2} 1079^{2} 2158^{2} 1079^{2}
(13, −125)  3660286792808760  639^{2} 1278^{2}  639^{2} 2556^{2}
(13, −69)  16837319246889384  516^{2}  516^{2}  258^{2} 1032^{2}
(13, −42)  20497606039673280  502^{2} 2008^{2}  251^{2} 2008^{2}
(13, −17)  12444975095505720  348^{2} 1392^{2}  348^{2} 2784^{2}
(13, −5)  3660286792794360 1583^{2} 1583^{2} 1583^{2} 3166^{2}
(13, −3)  1464114717117648 2364^{2} 2364^{2} 1182^{2} 2364^{2}
(13, 60)  457535849098320  552^{2} 1104^{2}  276^{2}  552^{2}
(13, 66)  32210523776515392  618^{2} 2472^{2}  309^{2} 2472^{2}
(13, 73)  610744996281840  494^{2} 1964^{2}  247^{2}  988^{2}
(13, 96)  10765549390536  588^{2} 1176^{2}  294^{2}  588^{2}
(13, 136)  264786704158368  258^{2} 1032^{2}  258^{2} 1032^{2}
(13, 544)  3111243773819208  929^{2} 1858^{2}  929^{2}  929^{2}
(13, 708)  21595692076981920  812^{2} 3248^{2}  406^{2} 1624^{2}
(13, 876)  835002924582096  340^{2} 1360^{2}  85^{2} 1360^{2}
(13, 928)  5307415849389480  470^{2} 1880^{2}  470^{2}  940^{2}
(14, −948)  2033174929441680  312^{2} 1248^{2}  156^{2}  624^{2}
(14, −800)  8235645283809960  390^{2} 1560^{2}  195^{2} 1560^{2}
(14, −672)  11529903397328568 2310^{2} 4620^{2} 2310^{2} 2310^{2}
(14, −596)  61355557364338608  598^{2} 1196^{2}  598^{2} 2392^{2}
(14, −281)  15300603799975032  253^{2} 1012^{2}  253^{2} 2024^{2}
(14, −212)  21824460002049648  560^{2}  560^{2}  560^{2} 1120^{2}
(14, −33)  72473678497325160 1002^{2} 2004^{2} 1002^{2} 4008^{2}
(14, −12)  3294258113514528 1077^{2} 2154^{2} 1077^{2} 2154^{2}
(14, −11)  144947356994638704  1806^{2} 3612^{2}  903^{2} 3612^{2}
(14, −3)  775119556121040  588^{2} 1176^{2}  294^{2} 1176^{2}
(14, 12)  205891132094640  564^{2} 2256^{2}  282^{2} 4512^{2}
(14, 96)  1647129056756616  306^{2} 1224^{2}  153^{2}  612^{2}
(14, 100)  16471290567565920 1186^{2} 2372^{2}  593^{2} 2372^{2}
(14, 240)  8235645283778760 1184^{2} 2368^{2}  592^{2} 1184^{2}
(14, 268)  726037150017264  858^{2} 1716^{2}  429^{2} 1716^{2}
(14, 528)  18118419624294264  356^{2} 1424^{2}  356^{2}  712^{2}
(14, 652)  33560254531348080  268^{2} 2144^{2}  67^{2} 2144^{2}

Table 2. Examples of elliptic curves E(n, p) (n = 14, 15; 0 < |p| 1000) with max_{1 i 4} |� _{i}| 1000^{2}.

(n, p)  N(n, p)  |� |  |� _{2}|  |� _{3}|  |� _{4}|
(15, −852)  8222777088032880  562^{2}  1124^{2}  281^{2}  1124^{2}
(15, −248)  141399694410862368  1185^{2}  4740^{2}  1185^{2}  4740^{2}
(15, −240)  74120807554080840  965^{2}  3860^{2}  965^{2}  3860^{2}
(15, −212)  280600200026160  498^{2}  1992^{2}  249^{2}  3984^{2}
(15, −116)  107475170953411824  2368^{2}  4736^{2}  2368^{2}  9472^{2}
(15, −96)  14824161510815304  1434^{2}  2838^{2}  717^{2}  2838^{2}
(15, −84)  3242785330490832  775^{2}  1650^{2}  775^{2}  1650^{2}
(15, −80)  74120807554076040  679^{2}  1358^{2}  679^{2}  1358^{2}
(15, −48)  14824161510815016  3057^{2}  3057^{2}  3057^{2}  3057^{2}
(15, −12)  5929664604325920  576^{2}  1152^{2}  288^{2}  1152^{2}
(15, −6)  237186584173036224  3705^{2}  3705^{2}  3705^{2}  7410^{2}
(15, −1)  59296646043258936  162^{2}  648^{2}  81^{2}  1296^{2}
(15, 1)  118593292086517776  4032^{2}  8064^{2}  2016^{2}  8064^{2}
(15, 12)  336912761609424  240^{2}  1920^{2}  60^{2}  3840^{2}
(15, 60)  37060403777035920  2299^{2}  4598^{2}  2299^{2}  2299^{2}
(15, 88)  130452621295164960  1232^{2}  2464^{2}  1232^{2}  2464^{2}
(15, 172)  2489995878769488  1258^{2}  2516^{2}  629^{2}  1258^{2}
(15, 375)  26953020928749960  1143^{2}  4572^{2}  1143^{2}  4572^{2}
(16, −408)  72579094756950240  1863^{2}  3726^{2}  3726^{2}  3726^{2}
(16, −96)  133417453597333128  3804^{2}  7608^{2}  1902^{2}  7608^{2}
(16, −33)  234814718331305640  3717^{2}  7437^{2}  3717^{2} 14868^{2}
(16, −32)  133417453597332744  5463^{2} 10926^{2}  5463^{2} 10926^{2}
(16, −8)  106733962877866080  891^{2}  891^{2}  891^{2}  1782^{2}
(16, 12)  2084647712458320  792^{2}  3168^{2}  396^{2}  6336^{2}
(16, 48)  7021971241964856  4608^{2}  9216^{2}  2304^{2}  9216^{2}
(16, 92)  61372028654772720  1064^{2}  2128^{2}  532^{2}  2128^{2}
(16, 268)  279342793469411664  2916^{2} 11664^{2}  1458^{2} 11664^{2}
(16, 300)  166771816996663440  1018^{2}  4072^{2}  509^{2}  4072^{2}
(16, 472)  186310763603371680  3119^{2} 12476^{2}  3119^{2} 12476^{2}
(16, 588)  116740271897662896  549^{2}  2196^{2}  549^{2}  1098^{2}
(16, 592)  17950711938549720  2221^{2}  8884^{2}  2221^{2}  4442^{2}
(16, 624)  102025111574427912  1100^{2}  2200^{2}  550^{2}  1100^{2}
(17, −404)  118434048164038608  3246^{2}  6492^{2}  1623^{2} 12948^{2}
(17, −68)  10206435200195943696  8284^{2} 33136^{2}  4142^{2} 33136^{2}
(19, −32)  19452264734491086120  31704^{2} 63408^{2} 31704^{2} 63408^{2}

Table 3. Examples of elliptic curves E(n, p) (16 n 19; 0 < |p| 1000) with max_{1 i 4} |� _{i}| 1000^{2}.

2 Values of the Goldfeld-Szpiro ratio GS(E)

The Goldfeld-Szpiro ratio was defined in (3). The articles of de Weger [We] and Nitaj [Ni] produce altogether 58 examples of elliptic curves with GS(E) greater than 1 (the record value being 42.265...). For all of these examples the conductor does not exceed 10^{10}. The largest values of GS(E) that we observed for our curves are tabulated in Table 4.

\begin{matrix}E & |� (E)| & GS(E) \\ E_{2}(9, 544) & 344^{2} & 1.20290... \\ E_{2,4} (16, 48) & 9216^{2} & 1.01357... \\ E_{2}(10, 204) & 504^{2} & 0.98366... \\ E_{4}(15, −212) & 3984^{2} & 0.94753... \\ E_{2,4} (19, −32) & 63408^{2} & 0.91159... \\ E_{4}(16, 12) & 6336^{2} & 0.87925... \\ E_{4}(15, 12) & 3840^{2} & 0.80334... \\ E_{2}(16, 592) & 8882^{2} & 0.58908... \\ E_{2}(11, 160) & 322^{2} & 0.57131... \\ E_{4}(17, −404) & 12984^{2} & 0.48986... \\ E_{4}(16, −33) & 14868^{2} & 0.45618... \\ E_{2}(13, 96) & 1176^{2} & 0.42149... \\ E_{2,4} (16, 472) & 12476^{2} & 0.36060... \\ E_{2,4} (17, −68) & 33136^{2} & 0.34368... \\ E_{2,4} (16, −32) & 10926^{2} & 0.32682... \\ E_{2}(11, 336) & 1058^{2} & 0.28146... \\ E_{4}(15, −116) & 9472^{2} & 0.27367... \\ E_{2,4} (16, 268) & 11664^{2} & 0.25741... \end{matrix}

Table 4. Elliptic curves E_{i}(n, p) (9 n 19; 0 < |p| 1000; 1 i 4) with the largest GS(E). Notation E_{i,j} (n, p) means that the given values of |� (E)| and GS(E) are shared by the isogeneous curves E_{i}(n, p) and E_{j} (n, p).

3 Large and small (nonzero) values of L(E, 1)

In this section we produce elliptic curves of rank zero with L(E, 1) either much smaller or much bigger than in all previously known examples (Tables 5 and

\begin{matrix}E & L(E, 1) \\ E(11, −733) & 88.203561907255071... \\ E(13, −160) & 71.523635814751843... \\ E(12, 466) & 56.224807584564927... \\ E(7, −433) & 36.275918867296195... \\ E(10, 687) & 30.274774697662334... \\ E(9, 767) & 29.638568367562609... \\ E(9, −93) & 28.032198538875886... \\ E(11, 336) & 22.922225180212583... \end{matrix}

Table 5. Elliptic curves E(n, p) (n 19; 0 < |p| 1000) with the largest values of L(E, 1) known to us.

\begin{matrix}E & L(E, 1) \\ E(12,800) & 0.0001706491750110... \\ E(10,142) & 0.0002457348122099... \\ E(11,168) & 0.0003276464160384... \\ E(14,672) & 0.0006067526222261... \\ E(9,160) & 0.0007372044423472... \\ E(10,-534) & 0.0009829392448696... \\ E(10,408) & 0.0009829392504019... \end{matrix}

Table 6. Elliptic curves E(n, p) (n 19; 0 < |p| 1000) with the smallest positive values of L(E, 1) known to us.

Note that

L(E(10, 408), 1) − L(E(10, −534), 1) = 0.00000000000553237117... .

This is the smallest known difference between the values of L(E, 1) of two elliptic curves of rank zero. The analytic orders of the Tate-Shafarevich group are 2^{2}, 4^{2}, 1^{2}, 4^{2} for the isogeneous curves E(10, 408) and E_{i}(10, 408), respectively, and 2^{2}, 8^{2}, 8^{2}, 8^{2} for the curves E(10, −534) and E_{i}(10, −534), repectively, where i = 2, 3, or 4.

We observed that for a large percentage of rank zero curves E_{i}(n, p) with 7 n 19 and 0 < |p| 1000, one has

L(E, 1) \frac{1}{(logN(E))^{2}} .

We verified, in particular, that (5) holds for every single curve E(7, p) of rank zero when 0 < |p| 1000. This is consistent with results of Iwaniec and Sarnak who proved [IS] that

L(f, 1) \frac{1}{(logN)^{2}} ,

for a large percentage of newforms of weight 2, with the level N of a newform f playing the role of the conductor of an elliptic curve.

On the other hand, we have

L(E(8, −131), 1) = 0.0002764516... < 0.0012048710... = (logN(8, −131))^{−2} ,

L(E(9, 160), 1) = 0.0007372044... < 0.0015186182... = (logN(9, 160))^{−2} ,

L(E(10, 142), 1) = 0.0002457384... < 0.0009026601... = (logN(10, 142))^{−2} ,

L(E(11, 168), 1) = 0.0003276464... < 0.0009902333... = (logN(11, 168))^{−2} ,

L(E(12, 800), 1) = 0.0001706491... < 0.0009613138... = (logN(12, 800))^{−2} .

An estimate much weaker than (5) was proposed by Hindry [H], see Conjecture 6 below.

4 Remarks on Conjecture 3

Below we sketch how to utilize curves E(n, p) in order to establish the first of the two conjectures of de Weger (Conjecture 3 above).

According to Chen [Ch], every sufficiently large even integer can be represented as the sum p + q where p is an odd prime and q is the product of at most two primes. Apply this, for sufficiently large n, to the number

4 · 3^{2n+1} = p + q.

The factors c_{∞} (E(n, p)) and c_{fin}(E(n, p)) on the right-hand-side of the formula for the analytic order of the Tate-Shafarevich group, (4), are given by the following lemma.

lemma Assume p < q, with q having at most two prime factors. Then we have

c_{∞} (E(n, p)) = \frac{π}{3^{n+1/2} · AGM(1, \sqrt{q/(p + q)})}  (4)

and

c_{fin}(E(n, p)) = 2c_{2}c_{3}c_{q}, i  (5)

(6)

where AGM(a, b) denotes the arithmetico-geometric mean of a and b,

c_{2} = 2 if p ≡ 1 mod 4 4 if p ≡ 3 mod 4 , c_{3} = 2(2n + 1) if p ≡ 2 mod 3 4 if p ≡ 1 mod 3 ,

and

c_{q} = \{ \begin{matrix}2 & if q is a prime \\ 4 & if q is a product of two primes \end{matrix} .

The conductor is given by the formula

N(E(n, p)) = 2^{f_{2}} · 3 · p · rad(q),

where rad(q) denotes the product of prime factors, and

f_{2} = \{ \begin{matrix}3 & if p ≡ 1(mod 4) \\ 4 & if p ≡ 3(mod 4) . \end{matrix}

lemma

This is easily proven by using calculations of Nitaj [Ni, Propositions 2.1, 3.1 and 3.2]. The following then seems to be a plausible conjecture.

Conjecture 5. For any ϵ > 0 there exists c(ϵ) > 0 and infinitely many n admitting a decomposition (6) with

p c(ϵ)q^{ϵ}

such that curve E(n, p) has rank zero.

If we accept Conjecture 5, then

\frac{1}{c_{∞} (E(n, p))} ≫ N(E(n, p))^{1/2−ϵ} and \frac{1}{c_{fin}(E(n, p))} ≫ N(E(n, p))^{−ϵ} .

on an infinite set of curves E(n, p).

Since |E(ℚ)_{tors} | 1 (in fact, |E(ℚ)_{tors} | can take only twelve values between 1 and 16, cf˙[Mz]) it remains to estimate L(E, 1). The result of Iwaniec and Sarnak mentioned in section 3 provides a support for the following conjecture recently proposed by Hindry [H, Conjecture 5.4].

Conjecture 6 (Hindry). One has

L^{(r)} (E, 1) ≫ N(E)^{−ϵ} (r being the rank of E).

Hindry observed that (8) implies that the distance from 1 to the nearest zero of L(E, s) is ≫ N(E)^{−ϵ} .

The combination of (7) and (8), for curves of rank zero, yields the assertion of Conjecture 3 for the analytic order of the Tate-Shafarevich group. In order to pass to the actual order, one needs, of course, the equality of the two, as predicted by the Birch and Swinnerton-Dyer Conjecture.

References

[BCDT] C. Breuil, B. Conrad, F. Diamond and R. Taylor, On the modularity of elliptic curves over ℚ: wild 3-adic exercises, J. Amer. Math. Soc. 14 (2001), 843–939

[Be] C. D. Beaver, 5-torsion in the Shafarevich-Tate group of a family of elliptic curves, J. Number Theory 82 (2000), 25–46

[C] J. W. S. Cassels, Arithmetic on curves of genus 1. VI. The Tate-Šafarevič group can be arbitrarily large, J. reine angew. Math. 214/215 (1964), 65–70

[Ch] J. R. Chen, On the representation of a large even number as the sum of a prime and the product of at most two primes, Sc. Sinica 16 (1973), 157–176

[Co] H. Cohen, A course in Computational Algebraic Number Theory. Graduate Texts in Math. 138 Springer-Verlag (1993)

[Cr_{1}] J. E. Cremona, Algorithms for modular elliptic curves. Cambridge University

Press 1997 [Cr_{2}] J. E. Cremona, Elliptic Curve Data (an online resource) [CFNS^{2}] J. E. Cremona, T. A. Fischer, C. O’Neil, D. Simon, M. Stoll, Ex-

plicit n-descent on elliptic curves. I. Algebra, II. Geometry, preprints 2006

[F] T. Fisher, Some examples of 5 and 7 descent for elliptic curves over ℚ, J. Eur. Math. Soc. 3 (2001), 169–201

[Fr_{1}] G. Frey, Der Rang der Lösungen von y^{2} = x^{3} ± p^{3} über ℚ, Manuscr. math.

48 (1984), 71–101 [Fr_{2}] G. Frey, A relation between the value of the L-series of the curve y^{2} = x^{3} −k^{3}

in s = 1 and its Selmer group, Arch. Math. 45 (1985), 232–238

[GHP] D. Goldfeld, J. Hoffstein, S. J. Patterson, On automorphic functions of half-integral weight with applications to elliptic curves. In: Number Theory Related to Fermat’s Last Theorem (Birkhäuser, Boston, MA, 1982), pp 153–193

[GL] D. Goldfeld and D. Lieman, Effective bounds on the size of the Tate-Shafarevich group, Math. Res. Letters 3 (1996), 309–318

[GS] D. Goldfeld and L. Szpiro, Bounds for the order of the Tate-Shafarevich group, Compos. math. 97 (1995), 71–87

[GA] C. Gonzalez-Avilés, On the conjecture of Birch and Swinnerton-Dyer, Trans. Amer. Math. Soc. 349 (1997), 4181-4200

[GJPST] G. Grigorov, A. Jorza, S. Patrikis, W.A. Stein, C. Tarniţă, Verification of the Birch and Swinnerton-Dyer conjecture for specific elliptic curves, Math. Comput. (to appear)

[H] M. Hindry, Why it is difficult to compute the Mordell-Weil group? In: Diophantine Geometry. Proceedings (ed. Umberto Zannier, Edizioni della Normale, Pisa, 2007)

[HL] J. Hoffstein, W. Luo, Nonvanishing of L-series and the combinatorial sieve, Math. Res. Letters 4 (1997), 435–442

[IS] H. Iwaniec and P. Sarnak, The non-vanishing of central values of automorphic L-functions and Landau-Siegel zeros, Israel J. Math. 120 (2000), 155–177

[K] V. A. Kolyvagin, Finiteness of E(ℚ) and � (E/ℚ) for a subclass of Weil curves, Izv. Akad. Nauk SSSR, 52 (1988), 522-540; translation in Math. USSR-Izv. 32 (1989), 523–541

[L] S. Lang, Conjectured diophantine estimates on elliptic curves. In: Arithmetic and Geometry – Papers dedicated to I. R. Shafarevich. Vol. 1 (Birkhäuser, Boston, MA, 1983), pp. 155–171 [Le] J. L. Lehman, Rational points on elliptic curves with complex multiplication by

the ring of integers in ℚ(\sqrt{−7} ), J. Number Theory 27 (1987), 253–272

[MM1] L. Mai and M. R. Murty, On the Shafarevich-Tate group. (unpublished, reported in [R])

[MM2] L. Mai and M. R. Murty, A note on quadratic twists of an elliptic curve. In: Elliptic Curves and Related Topics. CRM Proceed. and Lecture Notes 4 (1994), 121–124

[Ma] D. W. Masser, Note on a conjecture of Szpiro, Astérisque 183 (1990), 19–23 [Mz] B. Mazur, Modular curves and the Eisenstein ideal, Publ. Math. IHES 47

(1977), 33–186

[MS] J. R. Merriman, S. Siksek, N. P. Smart, Explicit 4-descents on an elliptic curve, Acta arithmetica 77 (1996), 385–404

[N] J. Nekovář, Class numbers of quadratic fields and Shimura’s correspondence, Math. Ann. 287 (1990), 577–594

[Ni] A. Nitaj, Invariants des courbes de Frey-Hellegouarch et grands groupes de Tate-Shafarevich, Acta arithmetica 93 (2000), 303–327

[O] K. Ono, Tate-Shafarevich groups of the congruent number elliptic curves, Acta arithmetica 81 (1997), 247–252

[R] C. S. Rajan, On the size of the Shafarevich-Tate group of elliptic curves over function fields, Compos. math. 105 (1997), 29–41

[Ro] D. E. Rohrlich, Unboundedness of the Tate-Shafarevich group in families of quadratic twists (an appendix to [HL]), Math. Res. Letters 4 (1997), 443–444

[Rs] H. E. Rose, On some elliptic curves with large Sha, Experimental Math. 9 (2000), 85–89

[Ru] K. Rubin, Tate-Shafarevich groups and L-functions of elliptic curves with complex multiplication. Invent. math 89 (1987), 527–559

[S] J. H. Silverman, The arithmetic of elliptic curves. Graduate Texts in Math. 106, Springer, New York 1986

[T] J. Tunnell, A classical diophantine problem and modular forms of weight 3/2, Invent. math. 72 (1983), 323–334

[W] J. L. Waldspurger, Sur les coefficients de Fourier des formes modulaires de poids demi-entiers, J. Math. Pure et Appl. 60 (1981), 375–484

[We] B. M. M. de Weger, A + B = C and big � ’s, Quart. J. Math. 49 (1998), 105–128

[ZK] D. Zagier and G. Kramarz, Numerical investigations related to the L-series of certain elliptic curves, Journal of the Indian Math. Soc. 52 (1987), 51–69